OpenStudy (anonymous):

need help with expressions 16^-3/4

OpenStudy (lgbasallote):

1/8

OpenStudy (anonymous):

how do you get to 1/8 ? steps would help thank you

OpenStudy (anonymous):

1) take the fourth root 2) cube the result 3) take the reciprocal

OpenStudy (lgbasallote):

^that

OpenStudy (anonymous):

\[16^{-\frac{3}{4}}=\frac{1}{\sqrt[4]{16}^3}=\frac{1}{2^3}=\frac{1}{8}\]

OpenStudy (anonymous):

thank you .. satellite .. how do you go about getting 1/2^3